# GUS-Net: Social Bias Classification in Text with Generalizations, Unfairness, and Stereotypes

### v2

@@ 5 Conclusion and Discussion
The proposed GUS-Net model addresses limitations in existing bias detection methods by focusing on the nuanced identification of social biases with semantic categories of generalizations, unfairness, and stereotypes. Moreover, GUS-Net uses a multi-label token classification architecture, based on bert-base-uncased, that allows entities to span multiple tokens and be nested within each other. GUS-Net approaches bias with three detailed entities, offering a more granular and precise detection of social biases. This enables better insights into the structural components of biased language. Our results demonstrate that GUS-Net performs well at classifying tokens as each of the entities, with a notable strength in detecting stereotypes. In sum, GUS-Net contributes the field of bias detection in NLP by incorporating a fine-grained and multi-faceted view of biased language.
Acknowledgements

### v3

@@ 7 Conclusion
The proposed GUS-Net model addresses limitations in existing bias detection methods by focusing on the nuanced identification of social biases with semantic categories of generalizations, unfairness, and stereotypes. We identify that auto-regressive decoder-only models are poorly suited for multi-label token classification. Conversely, BERT-base-uncased trained on the GUS dataset (GUS-Net) had superior alignment and effectiveness, while allowing for complex label structures that can span multiple words and can be nested/overlapping. GUS-Net approaches bias with three detailed entities, offering a more granular and precise detection of social biases. This enables better insights into the structural components of biased language. Our results demonstrate that GUS-Net performs well at classifying tokens as each of the entities, with a notable strength in detecting stereotypes. In sum, GUS-Net contributes the field of bias detection in NLP by incorporating a fine-grained and multi-faceted view of biased language.
Declaration
Authorship contribution statement
Maximus Powers: Project Lead, Engineering dataset generation, model training & evaluation, Writing drafting, review & editing. Shaina Raza: Advisor, Writing final draft, review & editing. Alex Chang: Writing review & editing, Dataset Quality human-in-the-loop review lead. Umang Mavani: Engineering synthetic corpus pipeline, Dataset Quality human-in-the-loop reviewer. Harshitha Reddy Jonala: Engineering annotation pipeline, Dataset Quality human-in-the-loop reviewer. Ansh Tiwari: Engineering annotation pipeline, Dataset Quality human-in-the-loop reviewer. Hua Wei: Advisor, Writing review & editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments